OpenStudy (vishweshshrimali5):

Solve the following equation in natural numbers: \(\large{x^2 + y^2 = 1997(x-y)}\)

OpenStudy (vishweshshrimali5):

@Abhishek619 @mathslover @Miracrown @mukushla @hartnn

OpenStudy (vishweshshrimali5):

@ganeshie8

OpenStudy (anonymous):

well, I think discriminant of quadratic must be a perfect square

hartnn (hartnn):

one obvious solution is 0,0 :P

mathslover (mathslover):

But that isn't Natural number..

OpenStudy (vishweshshrimali5):

Busted @hartnn

hartnn (hartnn):

and this is just a quadratic so you can use same method as for previous as mukushla said, discriminant = perfect square... yeah, ok ...find all integer solution and eliminate those are not applicable, that my way of solving :P

OpenStudy (vishweshshrimali5):

See how big my heart is !! Come on applaud and praise me.

Miracrown (miracrown):

I guess I would want to use a similar method to before. Express this as quadratic in "y" and then look for particular forms of the result that give us natural numbers. Let's try that Guess I am a stickler for basic methods

OpenStudy (lichking):

hmm

OpenStudy (vishweshshrimali5):

And I am going to stick with mine @Miracrown ;)

hartnn (hartnn):

170,145 ? 1827, 145 ?

Miracrown (miracrown):

sureeee do as you wish! :P

OpenStudy (dan815):

some obvious ones :) (0,0) , (1997,0), (0,-1997)

OpenStudy (vishweshshrimali5):

Yeah sure . Sorry the only answers are: (170,145); (1827,145); (1972,315)

OpenStudy (dan815):

oh nat numbers

mathslover (mathslover):

\(t = 1997 \sqrt{ 1 - \cfrac{2}{x+y} } \)

hartnn (hartnn):

you sure about 1972, 315 ?

OpenStudy (lichking):

i am not sure abt 1972

Miracrown (miracrown):

-4x^2 + 7988x + 1997^2 Okay, few notes We need this number to be a perfect square We also need it to be an odd perfect square because we need -1997 = odd... odd+odd = even so 2 divides evenly the root

OpenStudy (vishweshshrimali5):

Yeah sorry I am going through 2 questions at a time and mixing up answers.

OpenStudy (lichking):

yup i got the answer

OpenStudy (vishweshshrimali5):

The only solutions are (170,145); (1827,145)

OpenStudy (lichking):

1972 and 315 is correct

OpenStudy (lichking):

no my answeer is wrong

OpenStudy (lichking):

lemme check it again

OpenStudy (lichking):

i got 170,145 as my answer

Miracrown (miracrown):

So the only thing that I didn't write is that we almost must demand -4x^2+7988x + 1997^2 > 1997^2 because natural numbers are positive so This really limits where we are on the parabola

OpenStudy (vishweshshrimali5):

@Miracrown You are correct in saying that.

Miracrown (miracrown):

--> -4x^2 + 7988 x > 0 ---> 4x (-x + 1997) > 0. ---> 0 < x < 1997. Okay not that helpful >.<

Miracrown (miracrown):

Okay, the other constraints (-4x^2+7988x + 1997^2) = (2n+1) for some n (-4x^2+7988x+1997^2) = m^2 for some m

Miracrown (miracrown):

We can do quadratic formula on these again and get further conditions on x... but just thinking if that's not too messy

Miracrown (miracrown):

I am going to make some simplifications "b" = 1997 -4x^2 + 4 b x + b^2 -2n-1 = 0 x = -4b +- Sqrt[(4b)^2 - 4(-4)(b^2-2n-1)] all over -8

mathslover (mathslover):

Do the pythagoras substitution.

Miracrown (miracrown):

Likewise with m^2

Miracrown (miracrown):

What do you mean there? o.O you pythagorean triples

OpenStudy (anonymous):

well this is better, using some algebra\[(x+y)^2+(1997-x+y)^2=1997^2\] rest will be easy @mathslover ha?

Miracrown (miracrown):

x = -4b +- Sqrt[(4b)^2-4(-4)(b^2-m^2)] all over -8

mathslover (mathslover):

This is what I suggest : for : a^2 + b^2 = c^2 (m > n) a = (m^2 - n^2 ) b = 2mn c = m^2 + n^2

Miracrown (miracrown):

continuing with my work... x = b/2 -+ 2 Sqrt[b^2 - 2n - 1] Okay, this looks helpful. Because we want x to be a natural number. This means that we need b -+ 4 Sqrt[b^2-2n-1] to be even but b = 1997, which is odd odd+- even = odd which is never divisible by 2

Miracrown (miracrown):

this means we demand that b^2-2n-1 = 0

Miracrown (miracrown):

Ugh, this means n = 1994004

Miracrown (miracrown):

Actually, it's just a contradiction, because this means that x can never be a natural number. So this means there's something a bit wrong here. Rethinking

mathslover (mathslover):

a = x b = y c = \(\sqrt{1997(x-y)}\) a = m^2 - n^2 = x b = 2mn = y c = m^2 + n^2 = \(\sqrt{1997 (x-y)}\)

Miracrown (miracrown):

yea not sure how to work with that But I see the general idea of trying to fit the pythagorean triples to the answers

mathslover (mathslover):

http://www.mathblog.dk/pythagorean-triplets/ I am reading this blog right now. May be, any of you will also want to look at this.

Miracrown (miracrown):

This is just reminds me very much of diophantine type equations and fermat's last theorem type problems. But with the squares, you can relate it to the triples, which are pretty well known. So it irks me a bit not solving it completely. But yeah, I'll just get it later haha

mathslover (mathslover):

Yep you are right with diophantine type equations here. http://eprints.usq.edu.au/20783/1/Yevdokimov_ASMJ_v25n1_PV.pdf I have been doing some searches on google related to pythagorean triplet and found this. This will be solved by Pythagorean triplet, that I'm sure as we have to find another triplet satisfying the situation. But, am not sure how to do this.

OpenStudy (vishweshshrimali5):

Okay guys you did a lot of work now its my turn. I am going to continue from where @mukushla left; \((x+y)^2 + (1997-x+y)^2 = 1997^2\)

OpenStudy (vishweshshrimali5):

Now, since x and y are positive integers (or natural numbers), so, I can write that; \(0<x+y<1997\) and \(0<1997-x+y<1997\). Now I am putting \(a = x+y\) and \(b = 1997-x+y\). So, I only have to solve: \(a^2 + b^2 = 1997^2\) Thus, (a,b,1997) make Pythogorean triplets. Now, since, 1997 is a prime number. So, \(gcd(a,b) = 1\) . Now I am going to use Pythogorean substition, i.e. I am going to put, \(a = 2mn\) and \(b = m^2 - n^2\). Also, I am asuming m>n. And I know that gcd(a,b) = 1.

OpenStudy (vishweshshrimali5):

Now I am going to us modulo,

OpenStudy (vishweshshrimali5):

just a quick question how do I write this in latex ?

OpenStudy (rational):

\(\equiv\)

OpenStudy (vishweshshrimali5):

Thanks @rational. Now I have, \(m^2, n^2 \equiv 0,1,-1 (mod~5) ~\text{and}~ 1997 \equiv 2(mod~5)\) So, \(m,n \equiv \pm1~(mod~5)\) Similarly, Since, \(m^2, n^2 \equiv 0,1~ (mod~3) ~\text{and}~ 1997 \equiv 2~(mod~3)\) Thus, \(m,n = 1,4,11,14~(mod~15)\). Also, I assumen m>n. So, \[\cfrac{ 1997 }{ 2 } \le m^2 \le 1997\] Now \(\cfrac{1997}{2} = 998.5\) and \(\sqrt{998.5} =31.599\). Also, \(\sqrt{1997} = 44.68\). So, I only have to consider m = 34, 41, 44. But, only m = 34 satisfy the previous conditions. So, (m,n) = (34, 29) So, (a,b) = (1972, 315), which leads to our answer.

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 \(a = 2mn\) and \(b = m^2 - n^2\) (a and b can exchange themselves) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

Not the best solution but it works.

OpenStudy (vishweshshrimali5):

@ganeshie8 , @hartnn , @mukushla , @Miracrown

OpenStudy (vishweshshrimali5):

While many school students are not taught about modulo, it really helps in easily solving very difficult problems.